OpenStudy (jack1):

Would love a hand please: Solve the following system of partial differential equations for u(x,y): ∂u/∂y=2xyu ∂u/∂x=(y^2+5)u

OpenStudy (jack1):

@amistre64 @skullpatrol @waterineyes pretty please?

OpenStudy (jack1):

@terenzreignz any ideas anyone please? @Mertsj @dmezzullo or maybe @zzr0ck3r

OpenStudy (jack1):

so i know that integral of du = integral of 2*x*y*u dy... yeah...?

OpenStudy (jack1):

so \[u (y) = constant \times e ^{x y ^{2}}\] ... i think

OpenStudy (anonymous):

\[\frac{ \delta u }{ \delta y }=2xyu(x,y)\rightarrow \frac{ \delta u }{ u(x,y) }=2xy \delta y \rightarrow L(u(x,y))=xy^2+k(x)\]\[ L(u(x,y))=xy^2+k(x)\rightarrow u(x,y)=e^{k(x)+xy^2}=C(x)e^{xy^2}\]\[\frac{ \delta u(x,y) }{ \delta x}=\frac{ dC(x) }{ dx }e^{xy^2}+C(x)e^{xy^2}y^2=(y^2+5)C(x)e^{xy^2}\] Grouping terms:\[\frac{ dC(x) }{dx }=5\rightarrow L[C(x)]=5x \rightarrow C(x)=e^{5x}\]The solution is: \[u(x,y)=C(x)e^{xy^2}=e^{5x}e^{xy^2}=e^{xy^2+5x}\]

OpenStudy (anonymous):

Even better, the general solution is:\[u(x,y)=e^{xy^2+5x+c}\]

OpenStudy (jack1):

...whoa! dude that is awesome!! one quick Q, what is L in this equation...?

OpenStudy (anonymous):

Neperian Logarithm

OpenStudy (jack1):

that's brilliant @CarlosGP , cheers for that hey!

OpenStudy (anonymous):

Thanks, enjoy it!

OpenStudy (amistre64):

i would love to have the time to sit down and get a grasp on "systems of diffQs". we never got a chance to cover those in class

OpenStudy (anonymous):

@amistre64 We never had the time to cover this in class either, but this one seemed very intuitive